OpenStudy (anonymous):

Find the sum 11sqrt15 + 5sqrt15

OpenStudy (a_clan):

16sqrt15

OpenStudy (anonymous):

16sqrt15

OpenStudy (anonymous):

why r u asking these ....try taking the sqrt term common ..if u have so much difficulty on solving them directly

OpenStudy (anonymous):

Ayynadiax3 do know the answer to this: 11x +5x?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

16sqrt15

OpenStudy (amistre64):

the usual explanation is that when a person posts multiple times about the same exact topic; they are simply trying to get answers for the homework/test and want to move on to doing better things with their day :)

OpenStudy (moongazer):

you can add directly a radical if the radical is the same like \[11\sqrt{15} + 5\sqrt{15}\] just add 11 and 5 like what you do in 11x^2+5x^2 = 16x^2

OpenStudy (anonymous):

@ amistre64, then they should just learn to use wolframalpha.com! would save them even more time! haha